A fisherman notices that his boat is moving up and down periodically, owing to waves on the surface of the water. It takes 3.2 s for the boat to travel from its highest point to its lowest, a total distance of 0.58 m . The fisherman sees that the wave crests are spaced 4.2 m apart

Part A

How fast are the waves traveling?

Express your answer with the appropriate units.

Part B

What is the amplitude of each wave?

Express your answer with the appropriate units.

Part C

If the total vertical distance traveled by the boat were 0.35 m , but the other data remained the same, how fast are the waves traveling ?

Express your answer with the appropriate units.

Part D

If the total vertical distance traveled by the boat were 0.35 m , but the other data remained the same, what is the amplitude of each wave?

Express your answer with the appropriate units.
Answer #1

T = 2 * 3.2 = 6.4 s

L= 4.2

a)

velocity of wave

v = L / T = 4.2 / 6.4 = 0.6562 m/s

======

b)

A = 0.58 / 2 = 0.29 m

=====

c)

wave speed is independent of amolitude

v = 0.6562 m/s

======

d)

A = 0.35 / 2 = 0.175 m